OpenStudy (anonymous):

The square root of 238 is between which two numbers?

OpenStudy (unklerhaukus):

\[15^2=225\]\[x^2=238\]\[16^2=265\]

OpenStudy (anonymous):

\(16^2=256\) \(15^2=225\) Hence, \(15<\sqrt{238}<16 \)

OpenStudy (anonymous):

Thank you very much im not so good at math :)

OpenStudy (anonymous):

I am not either.

OpenStudy (unklerhaukus):

me either

OpenStudy (anonymous):

well i guess im not the smartest one out of you two lol